If cos x = - 1/3 is in third quadrant, then find sin x/2, cos x/2 and tan x/2


 


Grade:12th Pass

1 Answers

sanatan sharma
12 years ago

sin x/2 = √(2/3)

cos x/2 = 1/√3

tan x/2 = √2